A package contains 10 resistors, 2 of which are defective. If 3 are selected, find the probability of getting 1 defective resistor.
Found 2 solutions by Boreal, Alan3354:
Answer by Boreal(15235)

There are 3 ways to get 1 defective resistor, picked first, second, or third
8/10*7/9*2/8=112/720 and 3 ways so 336/720=0.46667
also (2C1)*8C2/10C3=0.46667

Answer by Alan3354(69443)

How is a resistor defective?
Does is fail to resist, and surrender?